# Dirac Equation Solution with Generalized tanh-Shaped Hyperbolic Potential: Application to Charmonium and Bottomonium Mass Spectra

### v1

@@ Acknowledgments
In the investigation of the quantum mechanical system described by Eq. (S44), the imposed boundary conditions stipulate that the wave functions $G_{nk}(0)=0$ and $G_{nk}(1)=0$, ensuring the physical realizability of the system within the defined spatial domain. Upon the application of a series of transformations, the system undergoes a significant reparameterization: the radial wave function $G_{nk}(r)$ transformed into $F_{nk}(r)$, the quantum number $k$ is inverted to $k+1$, the energy levels $E_{nk}$ are negated to $-E_{nk}$, and the constants and potentials are modified accordingly ($C_{ps}$rightarrow $-C_{s}$, $V_{1}$rightarrow $-V_{1}$, $V_{2}$rightarrow $-V_{2}$, $V_{3}$rightarrow $-V_{3}$). As a consequence of these transformations, Eq. (S44) and its associated expressions, denoted by Eq. (S45), evolve into Eqs. (S16) and (S17), respectively. This approach underscores the inherent symmetry between the pseudospin and spin symmetry frameworks, enabling a unified perspective on the energy spectrum and wavefunction solutions within the quantum mechanical paradigm.
@@
[…]
@@
The pseudospin symmetry plays a pivotal role in understanding the underlying principles governing the behavior of quantum systems. By delving into the energy eigenvalues equation, delineated by Eq. 4 of main part manuscript or Eq. (S49), under specific conditions, we elucidate the manifestation of this symmetry in diverse scenarios.
@@
From the above expression, when $k$=0 and $k$=1, we simplify it as:
@@
The transformation of pseudospin symmetry into spin symmetry can be achieved by applying the following series of transformations: $k$rightarrow $k+1$, $E_{nk}$rightarrow $-E_{nk}$, $C_{ps}$rightarrow $-C_{s}$, $V_{1}$rightarrow $-V_{1}$, $V_{2}$rightarrow $-V_{2}$, $V_{3}$rightarrow $-V_{3}$. This transformation emphasizes the intrinsic duality and interconnectedness of spin and pseudospin symmetries, offering a unified description of the energy spectrum across various quantum potentials. These findings not only deepen our understanding of pseudospin symmetry in quantum mechanics but also provide a pathway for experimentally testing theoretical predictions, thus bridging the gap between theory and experiment.
@@
where $E_{nl}$ is the energy eigenvalue obtained by using non-relativistic case Eq. 17 in main manuscript under the GTHP, and $\hslash$=$k$=$c$=1 are considered. For charmonium and bottomonium systems, the mass spectra can be expressed as:
@@
Using experimental data from Ref. [Particle Data Group et al., 2022], the potential parameters $V_{1}$, $V_{2}$, $V_{3}$, and $\alpha$ are determined as follows:
@@
This is because the ratio $\frac{\alpha}{\sqrt{2\mu}}$ is independent of quark mass. Consequently, $V_{2}\sim\left(\frac{\alpha}{\sqrt{2\mu}}\right)^{2}$ and $V_{3}\sim\left(\frac{\alpha}{\sqrt{2\mu}}\right)^{2}$ are also independent of quark mass, making the potential parameters for charm and bottom quarks the same. Note that for charm quarks,$\frac{\alpha}{\sqrt{2\mu}}=\frac{\alpha}{\sqrt{m_{c}}}=0.198424$ GeV1/2, and for bottom quarks, $\frac{\alpha}{\sqrt{2\mu}}=\frac{\alpha}{\sqrt{m_{b}}}=0.201649$ GeV1/2. Additionally, $V_{1}+m_{q}+m_{\bar{q}}=\text{const}$, implying $V_{1}+2m_{c}=9.9447$ GeV and $V_{1}+2m_{b}=13.2432$ GeV for charmonium and bottomonium, respectively. Hence, there is no difference between current and constituent quark masses for both charm and bottom quarks, justifying the use of $m_{c}=1.27$ GeV and $m_{b}=4.18$ GeV for mass spectrum calculations. The mass spectra of charmonium and bottomonium are computed and compared with experimental data as shown in Table S1 and S2.
@@
State Particle Present Work Present Work Present Work Theory I Theory II Theory III Experiment Non-Rel. Rel. $C_{s}$=0 [MeV] Rel. $C_{s}$=-62.1 [MeV] Ref. [Purohit et al., 2022] Ref. [Soni et al., 2018] Ref. [Bukor and Tekel, 2023] Ref.  [Particle Data Group et al., 2022] 1S J/$\psi$(1S) 3.096900 3.044003 3.096971 3.096 3.094 3.0969 3.09690 0.00% -1.71% $0$03% -0.09% 0.00% 1P $\chi_{c1}$(1P) 3.196126 3.108593 3.158390 3.415 3.468 3.518 3.51067 -8.96% -11.45% -10.03% -2.71% -1.22% 0.21% 1D $\psi$(3770) 3.377174 3.220174 3.265592 3.770 3.772 3.787 3.77370 -10.51% -14.67% -13.46% -0.10% -0.05% 0.35% 1F - 3.617957 3.359102 3.400419 4.040 4.012 - - 2S $\psi$(2S) 3.686100 3.496920 3.544020 3.733 3.681 3.686 3.68610 0.00% -5.13% -3.85% 1.27% -0.14% 0.00% 2P $\chi_{c1}$(3872) 3.772214 3.542490 3.588047 3.894 3.938 3.823 3.87165 -2.57% -8.50% -7.33% 0.58% 1.71% -1.26% 2D - 3.930322 3.625184 3.668331 4.088 4.188 - - 2F - 4.143411 3.733813 3.774369 - 4.396 - - 3S $\psi$(4040) 4.039000 3.806989 3.853836 4.068 4.129 3.889 4.03900 0.00% -5.74% -4.58% 0.72% 2.23% -3.71% 3P $\chi_{c1}$(4140) 4.106798 3.842917 3.888770 - 4.338 - 4.14650 -0.96% -7.32% -6.22% 4.62% 3D - 4.233365 3.909678 3.953877 - 4.557 - - 3F - 4.409120 4.000007 4.042271 - 4.746 - - 4S $\psi$(4230) 4.115000 4.011764 4.061630 4.263 4.514 3.982 4.22270 -2.55% -5.00% -3.81% 0.95% 6.90% -5.70% 4P $\chi_{c1}$(4274) 4.159579 4.041276 4.090428 - 4.696 - 4.28600 -2.95% -5.71% -4.56% 9.57% 4D - 4.246829 4.097048 $4$ 4.896 - -
@@
State Particle Present Work Present Work Present Work Theory I Theory II Theory III Experiment Non-Rel. Rel. $C_{s}$=0 [MeV] Rel. $C_{s}$=-36.9 [MeV] Ref. [Purohit et al., 2022] Ref. [Soni et al., 2018] Ref. [Bukor and Tekel, 2023] Ref. [Particle Data Group et al., 2022] 1S $\Upsilon$(1S) 9.460302 9.425395 9.460308 9.460 9.463 9.460 9.46030 $0$37% 0.00% 0.00% 0.03% 0.00% 1P $h_{b}$(1P) 9.594182 9.528470 9.562151 9.704 9.821 9.942 9.8993 -3.08% -3.75% -3.41% -1.97% -0.79% 0.43% 1D $\Upsilon_{2}$(1D) 9.827113 9.703603 9.735549 10.010 10.074 10.140 10.1637 -3.31% -4.53% -4.21% -1.51% -0.88% -0.23% 1F - 10.130188 9.921512 9.951663 10.268 10.288 - - 2S $\Upsilon$(2S) 10.023261 9.918959 9.951996 10.028 9.979 10.023 10.02326 0.00% -1.04% -0.71% 0.05% -0.44% 0.00% 2P $h_{b}$(2P) 10.151214 10.008571 10.040706 10.160 10.220 10.150 10.2598 -1.06% -2.45% -2.14% -0.97% -0.39% -1.07% 2D $\Upsilon_{2}$(2D) 10.380036 10.165760 10.196519 10.332 10.424 - - 2F - 10.688957 10.368270 10.397472 - 10.607 - - 3S $\Upsilon$(3S) 10.355200 10.246045 10.279213 10.343 10.359 10.178 10.3552 0.00% -1.05% -0.74% -0.12% 0.04% -1.71% 3P $h_{b}$(3P) 10.468567 10.324939 10.357396 - 10.556 - - 3D $\Upsilon_{2}$(3D) 10.681781 10.468012 10.499271 - 10.733 - - 3F - 10.986923 10.659500 10.689249 - 10.897 - - 4S $\Upsilon$(4S) 10.405400 10.409984 10.445219 10.536 10.683 10.242 10.5794 -1.64% -1.60% -1.27% -0.41% 0.98% -3.18% 4P $h_{b}$(4P) 10.497089 10.479373 10.514107 - 10.855 - - 4D $\Upsilon_{2}$(4D) 10.686643 10.611068 10.644635 - 11.015 - -
@@
The charmonium system, composed of a charm quark ($c$) and its anti-quark ($\bar{c}$), provides a fertile testing ground for understanding the interplay between relativistic effects and spin-orbit interactions. In the non-relativistic limit, the reasonably accurate description of the spectrum was observed, but discrepancies arise when compared with experimental results for higher orbital angular momentum states, such as $1P$, $1D$, and $1F$.
Upon introducing relativistic corrections to the mass spectra through the DE, while setting $C_{s}$ = 0 MeV (no spin symmetry breaking), significant deviations from experimental values are observed, especially for higher angular momentum states. For example, as shown in Table S1, the $\chi_{c1}(1P)$ state deviates by -11.45%, and the $\psi(3770)(1D)$ state shows an even larger deviation of -14.67%. These deviations are indicative of the inadequacies of a purely relativistic model when spin symmetry is assumed. Without spin symmetry breaking, the relativistic corrections tend to overestimate certain contributions, particularly in higher excited states, leading to noticeable discrepancies.
@@
To address these discrepancies, the spin symmetry breaking parameter $C_{s}$ was introduced, fine-tuned to ensure that the ground state $J/\psi(1S)$ matched the experimentally observed value. This adjustment allows for a closer comparison between theory and experiment, particularly for higher excited states. With $C_{s}=-62.1$ MeV the deviations in the $\chi_{c1}(1P)$ and $\psi(3770)(1D)$ states are reduced to -10.03% and -13.46% , respectively, as shown in Table S1.
While the introduction of $C_{s}$ improves the agreement with experimental data, it is evident that breaking spin symmetry alone is insufficient to fully reconcile theoretical predictions with experiment, particularly for the higher angular momentum states such as $2D$ and $3F$. The reason of this is that Eq. (S32) obtained spin symmetry case from radial DE. The persistent deviations suggest that additional spin-orbit interactions or higher-order corrections, such as tensor potentials, may be necessary. Furthermore, the DE overestimates certain corrections, indicating that a more sophisticated approach, such as solving a the two-body DE Crater et al. (2009); Scott et al. (1992); Ferreira (1988) and the Bethe-Salpeter equation Hilger et al. (2015); Soni et al. (2018), describing the structure of a relativistic two-body (particles) bound state, may provide a more accurate description of the charmonium system.
@@
The results from both the charmonium and bottomonium systems highlight the critical role of spin symmetry in determining higher exciting states in quarkonium systems. In the absence of spin symmetry breaking ($C_{s}$ = 0), the relativistic corrections cannot well to capture the experimentally observed fine structure, particularly for higher angular momentum states. This evident becomes especially evident in the charmonium system, where the fine structure splitting of states with higher orbital angular momentum, such as $1D$ and $1F$, cannot be captured without explicitly breaking spin symmetry. By introducing a non-zero $C_{s}$, we effectively break spin symmetry and improve the agreement between theoretical predictions and experimental data. However, the deviations that persist, particularly for higher excited states, suggest that breaking spin symmetry alone is not sufficient because Eq. (S32) obtained spin symmetry case from radial DE. To achieve a more accurate description of these states and also the experimentally observed splitting patterns, additional corrections, such as improved potential models incorporating tensor potentials or more sophisticated approaches like the two-body DE Crater et al. (2009); Scott et al. (1992); Ferreira (1988) and the Bethe-Salpeter equation Hilger et al. (2015); Soni et al. (2018), are required.

### v4

@@ Acknowledgments
In the investigation of the quantum mechanical system described by Eq. (S44), the imposed boundary conditions stipulate that the wave functions $G_{nk}(0)=0$ and $G_{nk}(1)=0$, ensuring the physical realizability of the system within the defined spatial domain. Upon the application of a series of transformations, the system undergoes a significant reparameterization: the radial wave function $G_{nk}(r)$ transformed into $F_{nk}(r)$, the quantum number $k$ is inverted to k-k, the energy levels $E_{nk}$ are negated to $E_{nk}$, and the constants and potentials are modified accordingly ($C_{ps}$ $\rightarrow$ $C_{s}$, $V_{1}$ $\rightarrow$ $V_{1}$, $V_{2}$ $\rightarrow$ $V_{2}$, $V_{3}$ $\rightarrow$ $V_{3}$). As a consequence of these transformations, Eq. (S44) and its associated expressions, denoted by Eq. (S45), evolve into Eqs. (S16) and (S17), respectively. This approach underscores the inherent symmetry between the pseudospin and spin symmetry frameworks, enabling a unified perspective on the energy spectrum and wavefunction solutions within the quantum mechanical paradigm.
@@
$\int^{\infty}_{0}{{\left|G_{nk}\left(r\right)\right|}^{2}dr}=\frac{1}{\alpha}\int^{1}_{0}{\frac{{\left|G_{nk}\left(z\right)\right|}^{2}dz}{1-z^{2}}}=1\ \ .$ (S51)
@@
The pseudospin symmetry plays a pivotal role in understanding the underlying principles governing the behavior of quantum systems. The analytic treatment under pseudospin symmetry can be also useful in nuclear physics contexts, since many nuclear single-particle levels exhibit pseudospin partners. Our results also suggest the GTHP could accommodate such features, hinting at applications in nuclear shell model studies. By delving into the energy eigenvalues equation, delineated by Eq. 4 of main part manuscript or Eq. (S49), under specific conditions, we elucidate the manifestation of this symmetry in diverse scenarios.
@@
From the above expression, when kk=0 and kk=1, we simplify it as:
@@
The transformation of pseudospin symmetry into spin symmetry can be achieved by applying the following series of transformations: $k$ $\rightarrow$ k-k, $E_{nk}$ $\rightarrow$ $E_{nk}$, $C_{ps}$ $\rightarrow$ $C_{s}$, $V_{1}$ $\rightarrow$ $V_{1}$, $V_{2}$ $\rightarrow$ $V_{2}$, $V_{3}$ $\rightarrow$ $V_{3}$. This transformation emphasizes the intrinsic duality and interconnectedness of spin and pseudospin symmetries, offering a unified description of the energy spectrum across various quantum potentials. These findings not only deepen our understanding of pseudospin symmetry in quantum mechanics but also provide a pathway for experimentally testing theoretical predictions, thus bridging the gap between theory and experiment.
@@
[…]
@@
Using experimental data from Ref. [75], the potential parameters $V_{1}$, $V_{2}$, $V_{3}$, and $\alpha$ are determined as follows:
@@
[…]
@@
[…]
@@
State Particle Present Work Present Work Present Work Theory I Theory II Theory III Experiment Non-Rel. Rel. $C_{s}$=0 [MeV] Rel. $C_{s}$=-36.9 [MeV] Ref. [61] Ref. [62] Ref. [63] Ref. [75] 1S $\Upsilon$(1S) 9.460302 9.425395 9.460308 9.460 9.463 9.460 9.46030 0.00% -0.37% 0.00% 0.00% 0.03% 0.00% 1P $h_{b}$(1P) 9.594182 9.528470 9.562151 9.704 9.821 9.942 9.8993 -3.08% -3.75% -3.41% -1.97% -0.79% 0.43% 1D $\Upsilon_{2}$(1D) 9.827113 9.703603 9.735549 10.010 10.074 10.140 10.1637 -3.31% -4.53% -4.21% -1.51% -0.88% -0.23% 1F - 10.130188 9.921512 9.951663 10.268 10.288 - - 2S $\Upsilon$(2S) 10.023261 9.918959 9.951996 10.028 9.979 10.023 10.02326 0.00% -1.04% -0.71% 0.05% -0.44% 0.00% 2P $h_{b}$(2P) 10.151214 10.008571 10.040706 10.160 10.220 10.150 10.2598 -1.06% -2.45% -2.14% -0.97% -0.39% -1.07% 2D $\Upsilon_{2}$(2D) 10.380036 10.165760 10.196519 10.332 10.424 - - 2F - 10.688957 10.368270 10.397472 - 10.607 - - 3S $\Upsilon$(3S) 10.355200 10.246045 10.279213 10.343 10.359 10.178 10.3552 0.00% -1.05% -0.74% -0.12% 0.04% -1.71% 3P $h_{b}$(3P) 10.468567 10.324939 10.357396 - 10.556 - - 3D $\Upsilon_{2}$(3D) 10.681781 10.468012 10.499271 - 10.733 - - 3F - 10.986923 10.659500 10.689249 - 10.897 - - 4S $\Upsilon$(4S) 10.405400 10.409984 10.445219 10.536 10.683 10.242 10.5794 -1.64% -1.60% -1.27% -0.41% 0.98% -3.18% 4P $h_{b}$(4P) 10.497089 10.479373 10.514107 - 10.855 - - 4D $\Upsilon_{2}$(4D) 10.686643 10.611068 10.644635 - 11.015 - -
@@
The charmonium system, composed of a charm quark ($c$) and its anti-quark ($\bar{c}$), provides a fertile testing ground for understanding the interplay between relativistic effects and spin-orbit interactions. In the non-relativistic limit, the reasonably accurate description of the spectrum was observed, but discrepancies arise when compared with experimental results for higher orbital angular momentum states, such as $1P$, $2P$, $3P$ and $4P$.
[…]
@@
[…]
While the introduction of $C_{s}$ improves the agreement with experimental data, it is evident that breaking spin symmetry alone is insufficient to fully reconcile theoretical predictions with experiment, particularly for the higher angular momentum states such as $\chi_{c1}(1P\longmapsto 4P)$. The reason of this is that Eq. (S32) obtained spin symmetry case from radial DE. The persistent deviations suggest that additional spin-orbit interactions or higher-order corrections, such as tensor potentials, may be necessary. Furthermore, the DE overestimates certain corrections, indicating that a more sophisticated approach, such as solving a the two-body DE [90, 91, 92] and the Bethe-Salpeter equation [93, 62], describing the structure of a relativistic two-body (particles) bound state, may provide a more accurate description of the charmonium system.
@@
The results from both the charmonium and bottomonium systems highlight the critical role of spin symmetry in determining higher exciting states in quarkonium systems. In the absence of spin symmetry breaking ($C_{s}$ = 0), the relativistic corrections cannot well to capture the experimentally observed fine structure, particularly for higher angular momentum states. This evident becomes especially evident in the charmonium system, where the fine structure splitting of states with higher orbital angular momentum, such as $\chi_{c1}(1P\longmapsto 4P)$, cannot be captured without explicitly breaking spin symmetry. By introducing a non-zero $C_{s}$, we effectively break spin symmetry and improve the agreement between theoretical predictions and experimental data. However, the deviations that persist, particularly for higher excited states, suggest that breaking spin symmetry alone is not sufficient because Eq. (S32) obtained spin symmetry case from radial DE. To achieve a more accurate description of these states and also the experimentally observed splitting patterns, additional corrections, such as improved potential models incorporating tensor potentials or more sophisticated approaches like the two-body DE [90, 91, 92] and the Bethe-Salpeter equation [93, 62], are required.
Notes and references